OpenStudy (anonymous):

Help!!!! Solve for x. x/14=4/7

OpenStudy (anonymous):

4*14 divided by 7 will give you x

OpenStudy (anonymous):

i don't get how to get x. Should i divide? multiply?

OpenStudy (anonymous):

oh okay. thank you!

OpenStudy (anonymous):

multiply both sides by 14

OpenStudy (anonymous):

so it would be x time 14= 4/7 times 14??

OpenStudy (anonymous):

no you have \[\frac{x}{14}\] so \[\frac{x}{14}\times 14=x\] on the left, then \[\frac{4}{7}\times 14\] on the right

OpenStudy (anonymous):

or maybe this is easier: \[\frac{x}{14}=\frac{4}{7}=\frac{8}{14}\implies x = 8\]

OpenStudy (anonymous):

okay! thanks guys!